OpenStudy (anonymous):

any idea how to solve this? 3. Bricks are delivered to a work site and stacked in rows and columns, forming a rectangular prism. The length of the prism is 1 foot greater than its width, and its height is 2 feet less than its width. Find the dimensions of the prism formed by the bricks, given that its volume is 40 cubic feet.

OpenStudy (anonymous):

40=(length)x(width)x(height)

OpenStudy (anonymous):

length = width+1 height = width-2

OpenStudy (anonymous):

(w+1) (w) (w - 2) = 40

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

solve for width

OpenStudy (anonymous):

then you can find the other dimensions

OpenStudy (anonymous):

(w^2+1) (W - 2) = 40 would i multiply them then go from there? : W^3 - 2w^2 +1w -2 = 40

OpenStudy (anonymous):

40 = (w+1)(w)(w-2) = (w^2 +w)(w-2) = (w^3 -2w^2 +w^2 -2w) = (w^3 -w^2 -2w)

OpenStudy (anonymous):

then go from there.

OpenStudy (anonymous):

find the real solution of w

OpenStudy (anonymous):

what happened to 40? and would i use synt. div. to find w?

OpenStudy (anonymous):

40 is still there

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

0 = w^3 - w^2 -2w -40 factor it: 0 = (w-4)(w^2 - 3w +10) then you get your real solutions of w = 4 or you find the real solution of (w^2 - 3w +10) but since we already have w = 4. thats the answer

OpenStudy (anonymous):

you can split 0 = (w-4)(w^2 - 3w +10) into 2 equations 0 = w-4 and 0 = (w^2 - 3w +10)

OpenStudy (anonymous):

0 = w-4 4 = w 0 = w^2 - 3w + 10 -10 = w^2 -3w -3.3 = w is this correct?

OpenStudy (anonymous):

well you dont need the second equation. you already have 4. and there wouldnt be -dimensions in the question. so your width = 4

OpenStudy (anonymous):

now you can find your length and height from the width

OpenStudy (anonymous):

oh i see so (4+1) x 4 x (4-2) = 40?

OpenStudy (anonymous):

you tell me

OpenStudy (anonymous):

it does thank you :) so i had to factor it after getting w+1 x w x w - 2 = 40 :)

OpenStudy (anonymous):

i have to go lay down i cant stop coughing thank you again

OpenStudy (anonymous):

no problem